# Data-Driven Spectral Analysis Through Pseudo-Resolvent Koopman Operator in Dynamical Systems

### v1

@@ 4 Convergence Analysis of Spectrum
For any $f\in\mathcal{F}$, we have
$\mathcal{P}_{N}f-\mathcal{P}f=\frac{1}{2\pi i}\int_{\Gamma}[\mathcal{R}_{N}(z)-\mathcal{R}(z)]x\,dz.$
Using the resolvent identity, whenever $N$ is large enough we get
[…]
$\leq\frac{|\Gamma|}{2\pi}\sup_{z\in\Gamma}\|\mathcal{R}_{N}(z)\|\cdot\sup_{z\in\Gamma}\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{R}(z)x\|,$
Thus, by Lemma 4.7, we have $\text{dim}\,\mathcal{P}_{N}\mathcal{F}\geq\text{dim}\,\mathcal{P}\mathcal{F}$, for $N$ large enough. ∎
@@
in the Kuratowski sense [chaitin1983spectral, Chapter 5.1.1] (hence, implying Hausdorff convergence). Moreover, if Assumption 4.5 holds, we have $\sigma(\mathcal{K}_{N})\cap\Delta$ consists of exactly $m$ eigenvalues, for $N$ large enough.
@@
We will first show the lower semicontinuity and upper semicontinuity (See Remark  4.11), which implies the Hausdorff convergence, then we will show that the algebraic multiplicity is preserved. Let $\mathcal{P}$ and $\mathcal{P}_{N}$ be the spectral projections of $\mathcal{K}$ and $\mathcal{K}_{N}$ associated with $\lambda$ and $\sigma(\mathcal{K}_{N})\cap\Delta$, respectively.
Lower Semicontinuity: Choose $\varepsilon>0$ small enough so that $\Delta_{\varepsilon}=\Delta-\{z:|z-\lambda|<\varepsilon\}$ and the circle $\gamma_{\varepsilon}=\{z:|z-\lambda|=\varepsilon\}$ are in $\rho(\mathcal{K})$. Stability from Lemma 4.6 ensures a bound $\sup_{z\in\Delta_{\varepsilon}}\|\mathcal{R}_{N}(z)\|\leq M(\varepsilon)$ for $N>N(\varepsilon)$. By Corollary 4.9, $\dim\mathcal{P}^{\gamma_{\varepsilon}}_{N}\mathcal{F}\geq m$ for $N>N(\varepsilon)$ implies $\sigma(\mathcal{K}_{N})\cap\{z:|z-\lambda|<\varepsilon\}\neq\emptyset$. Thus, there exists $\lambda_{N}\in\sigma(\mathcal{K}_{N})\cap\Delta$ with $|\lambda_{N}-\lambda|<\varepsilon$ for large $N$, confirming $\lambda$ as a limit point.
Upper Semicontinuity: Lemma 4.6 ensures $\Delta_{\varepsilon}\subset\rho(\mathcal{K}_{N})$ for $N>N(\varepsilon)$, so $\sigma(\mathcal{K}_{N})\cap\Delta_{\varepsilon}=\emptyset$. Hence, any $\mu\in\sigma(\mathcal{K}_{N})\cap\Delta$ satisfies $|\mu-\lambda|<\varepsilon$. As $\varepsilon\to 0$, any subsequence of $\mu$ converges to $\lambda$.
Multiplicity Preservation: Suppose Assumption 4.5 holds. Now we show that $\sigma(\mathcal{K}_{N})\cap\Delta$ consists of exactly $m$ eigenvalues, counting algebraic multiplicities. By [chaitin1983spectral, Theorem 2.27], $\sigma(\mathcal{K}_{N})\cap\Delta$ is the spectrum of $\mathcal{P}_{N}\mathcal{K}_{N}\mathcal{P}_{N}$ on $\mathcal{P}_{N}\mathcal{F}$. Since $\dim\mathcal{P}_{N}\mathcal{F}=m<\infty$, $\mathcal{P}_{N}\mathcal{K}_{N}\mathcal{P}_{N}$ is a finite-dimensional operator with at most $m$ eigenvalues (counting multiplicities). If there were more, $\dim\mathcal{P}_{N}\mathcal{F}>m$, contradicting the assumption; if fewer, $\dim\mathcal{P}_{N}\mathcal{F}<m$, also a contradiction for large $n$. Thus, $\sigma(\mathcal{K}_{N})\cap\Delta$ consists of exactly $m$ eigenvalues, which preserves the algebraic multiplicity. ∎
@@
Intuitively speaking, upper semicontinuity ensures that every limit of approximate spectral points is contained in the true spectrum, which prevents irrelevant points (i.e., spectral pollution) and lower semicontinuity ensures that every point of the true spectrum is approached by some approximate spectral points, which prevents loss of spectral points (i.e., spectral inlusion).
@@
The [chaitin1983spectral, Theorem 2.27] in the above theorem mainly says that one can construct a spectral projection via a contour integral, which decomposes the space into two invariant subspaces: one that contains exactly the isolated part of the spectrum and one that contains the remainder. Essentially, it provides a method to “split” or reduce $\mathcal{K}$ into simpler parts corresponding to these separated spectral regions.
@@
Now we want to extend the Hausdorff convergence result beyond isolated eigenvalues to a compact subset of the spectrum. In order to do so, we need to introduce a new concept called resolvent stability as following.
@@
Let $\mathcal{K}$ be a bounded Koopman operator and $\{\mathcal{K}_{N}\}_{N\in\mathbb{N}}$ be finite-dimensional approximations that satisfy Assumption 4.2. Let $U\subset\sigma(\mathcal{K})$ be a non-empty compact subset enclosed by a positively oriented contour $\Gamma\subset\rho(\mathcal{K})$, and let $\Delta$ denote the interior of $\Gamma$. Then
@@
Because $\Delta\setminus U\subset\rho(\mathcal{K})$, every $z\in\Delta\setminus U$ lies in the resolvent set of $\mathcal{K}$. Assumption 4.2 therefore yields constants $N_{z}\in\mathbb{N}$ and $M_{z}>0$ such that for all $N>N_{z}$,
@@
where $\mathcal{R}_{N}(z)=(zI-\mathcal{K}_{N})^{-1}$. Fix $z\in\Delta\setminus U$ and assume $\liminf_{N\to\infty}\operatorname{dist}\bigl(\sigma(\mathcal{K}_{N}),z\bigr)>0.$ Then there exist $\delta>0$ and $N_{0}$ such that $\operatorname{dist}(\sigma(\mathcal{K}_{N}),z)>\delta$ for all $N>N_{0}$. Consequently, for every $N>\max\{N_{0},N_{z}\}$ we have the uniform bound $\|\mathcal{R}_{N}(z)\|\leq M_{z}$. Hence the sequence $\{\mathcal{K}_{N}\}$ is resolvent-stable around $U$ in the sense of Definition 4.14. Applying Mills [10.1137/0716052, Theorem 3.2] gives $\liminf_{N\to\infty}\operatorname{dist}\bigl(U,\sigma(\mathcal{K}_{N})\cap\Delta\bigr)\to 0$, which is exactly the desired lower semicontinuity. ∎
@@
Upper semicontinuity in Theorem 4.16 is not satisfied unless we have a stronger condition that $\mathcal{K}_{N}$ uniformly converges to $\mathcal{K}$. Refer to [kato1995perturbation, Theorem 3.1, Chapter 4] for more details.
We want to point out that Chatelin’s work in [chaitin1983spectral] primarily addresses the convergence of isolated eigenvalues which provides a comprehensive framework that ensures preservation of algebraic multiplicities during the approximation process. This is particularly valuable for numerical analysis as it guarantees that the number of approximate eigenvalues (counting multiplicity) matches that of the exact operator. In contrast, Mills’ work in [10.1137/0716052] extends spectral convergence theory (i.e., lower semicontinuity) to compact subsets of the spectrum, which offers a much broader applicability beyond just isolated eigenvalues.
@@ 5 Error Bounds Analysis
In last section, we established the convergence of the approximated spectrum $\sigma(\mathcal{K}_{N})$ to the true spectrum $\sigma(\mathcal{K})$. However, for practical application, it is essential to quantify the error of a computed solution. In this section, we will derive an error bound for the computed eigenvalues that is controlled by a computable residual quantity $\eta_{N}$, as defined later. This provides a practical way to assess the accuracy of the spectral information obtained via Galerkin methods such as EDMD [Williams_2015] and our pseudo-resolvent approach.
@@
Let $\mathcal{P}_{N}^{\lambda}$ be the spectral projection associated with the eigenvalues of $\mathcal{K}_{N}$ inside the contour $\Gamma_{\lambda}$. We define the computable residual quantity $\eta_{N}$ as:
$\eta_{N}:=\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|.$ (10)
To support our main proof, we state a key lemma from spectral approximation theory in the following which establishes the fundamental relationship between the error in the mean eigenvalue and the trace of an operator difference.
Lemma 5.3 ([chaitin1983spectral, Lemma 6.2]).
@@
Let $\lambda$ be an isolated eigenvalue of the Koopman operator $\mathcal{K}$ with finite algebraic multiplicity $m_{\lambda}$. Let $\Gamma_{\lambda}$ be a closed Jordan curve isolating $\lambda$ from the rest of $\sigma(\mathcal{K})$. Suppose that Assumption 4.2 and Assumption 4.5 hold, then for $N$ sufficiently large, the error in the arithmetic mean of the approximate eigenvalues is bounded by the computable residual $\eta_{N}$, as given in the Definition 5.2. Specifically, there exists a constant $C$ independent of $N$ such that:
$|\lambda-\bar{\lambda}_{N}|\leq C\cdot\eta_{N}.$
@@
Let $\mathcal{M}_{N}^{\lambda}\coloneqq\mathcal{P}_{N}^{\lambda}\mathcal{F}$ be the $m_{\lambda}$-dimensional approximate invariant subspace. Define the operators $\mathcal{G}_{N}:=\mathcal{K}_{N}|_{\mathcal{M}_{N}^{\lambda}}$ and $\mathcal{H}_{N}:=\mathcal{P}_{N}^{\lambda}\mathcal{K}\mathcal{P}_{N}^{\lambda}$. The arithmetic mean $\bar{\lambda}_{N}$ satisfies $\bar{\lambda}_{N}=\frac{1}{m_{\lambda}}\operatorname{tr}(\mathcal{G}_{N})$ and the error bound derives from the trace difference satisfies $|\operatorname{tr}(\mathcal{H}_{N}-\mathcal{G}_{N})|\leq m_{\lambda}\cdot\|\mathcal{P}_{N}^{\lambda}(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|$ by Lemma 5.3. Then, by uniform boundedness of $\|\mathcal{P}_{N}^{\lambda}\|\leq C$ (from Assumption 4.5) and [chaitin1983spectral, Theorem 6.15], we obtain the following result:
$|\lambda-\bar{\lambda}_{N}|\leq C\cdot\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|=C\cdot\eta_{N}.$
@@ 6.1 Pendulum
In this experiment, data is generated by simulating a single pendulum system. In the simulation setting, a single trajectory of 1500 time steps is produced with a fixed sampling time step of 0.2 seconds. The simulation starts from a randomly initialized state where the angle $\theta$ is chosen uniformly from 0 to $2\pi$ and its angular velocity $\dot{\theta}$ from 5-5 to $5$. The system’s dynamics are governed by the equation
$\ddot{\theta}=-\frac{3g}{2l}\sin(\theta)$
@@
with the parameters set as follows: gravitational acceleration $g=9.8$ and pendulum length $l=1.0$. At each time step, the angular acceleration is calculated using this equation, and then the angular velocity and angle are updated via Euler’s method. To select the dictionary, we use two sets of basis functions: a Fourier basis for the angle $\theta$ and a Hermite polynomial basis for the angular velocity $\dot{\theta}$. Specifically, the Fourier basis is constructed using complex exponentials $\exp\{in\theta\}$ for $n$ ranging from 20-20 to $20$, and the Hermite basis employs Hermite polynomials evaluated up to order $30$ (with the velocity $\dot{\theta}$ rescaled by $1/\sqrt{2}$). These basis functions are multiplied to form a dictionary of observables, which is then used to approximate the resolvent of the Koopman operator and study the system’s spectral properties.
@@
In Figure 1, we illustrate the behavior of $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$ as the grid point $z$ varies across these circles with radii $\Gamma_{1.01}$, $\Gamma_{1.1}$, and $\Gamma_{1.5}$. Theoretically, the inverse of the spectral norm of the resolvent should approach zero as $z$ gets closer to the unit circle, where the true spectrum of the Koopman operator for this pendulum system is expected to locate. This is observed in the figure that the value of $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$ decreases as the radius approaches 1. Consequently, by setting a smaller threshold for detection, only grid points nearer to the unit circle are identified as part of the spectrum, which enhances the accuracy of the spectral approximation.

### v2

@@ 4 Convergence Analysis of Spectrum
Fix $f\in\mathcal{F}$. By Definition 4.4,
$(\mathcal{P}_{N}^{\lambda}-\mathcal{P}^{\lambda})f=\frac{1}{2\pi i}\int_{\Gamma_{\lambda}}\bigl(\mathcal{R}_{N}(z)-\mathcal{R}(z)\bigr)f\,dz.$
For each $z\in\Gamma_{\lambda}$, Eq. (8) implies $\mathcal{R}_{N}(z)f\to\mathcal{R}(z)f$ as $N\to\infty$. Moreover, Lemma 4.6 (applied to $U=\Gamma_{\lambda}$) yields constants $N_{\Gamma}$ and $M_{\Gamma}$ such that
$\sup_{z\in\Gamma_{\lambda},\,N>N_{\Gamma}}\|\mathcal{R}_{N}(z)\|\leq M_{\Gamma}.$
Hence, for any $z,w\in\Gamma_{\lambda}$ and $N>N_{\Gamma}$, the first resolvent identity gives
$\|\mathcal{R}_{N}(z)f-\mathcal{R}_{N}(w)f\|\leq|z-w|\,\|\mathcal{R}_{N}(z)\|\,\|\mathcal{R}_{N}(w)\|\,\|f\|\leq|z-w|\,M_{\Gamma}^{2}\,\|f\|.$
Thus $\{z\mapsto\mathcal{R}_{N}(z)f\}_{N>N_{\Gamma}}$ is equicontinuous and uniformly bounded on the compact set $\Gamma_{\lambda}$. Since the pointwise limit $z\mapsto\mathcal{R}(z)f$ is continuous, an Arzelà–Ascoli argument [6] implies the convergence is uniform on $\Gamma_{\lambda}$, i.e.,
$\sup_{z\in\Gamma_{\lambda}}\|(\mathcal{R}_{N}(z)-\mathcal{R}(z))f\|\to 0.$
Therefore,
$\|(\mathcal{P}_{N}^{\lambda}-\mathcal{P}^{\lambda})f\|\leq\frac{|\Gamma_{\lambda}|}{2\pi}\sup_{z\in\Gamma_{\lambda}}\|(\mathcal{R}_{N}(z)-\mathcal{R}(z))f\|\to 0,$
so $\mathcal{P}_{N}^{\lambda}\to\mathcal{P}^{\lambda}$ pointwise. Applying Lemma 4.7 yields $\dim(\mathcal{P}_{N}^{\lambda}\mathcal{F})\geq\dim(\mathcal{P}^{\lambda}\mathcal{F})$ for $N$ large enough. ∎
@@
in the Kuratowski sense [4, Chapter 5.1.1] (hence, implying Hausdorff convergence). Moreover, if Assumption 4.5 holds, we have $\sigma(\mathcal{K}_{N})\cap\Delta$ consists of exactly $m$ eigenvalues, for $N$ large enough.
@@
We show lower semicontinuity and upper semicontinuity (Remark 4.11), then prove multiplicity preservation.
Lower semicontinuity. Choose $\varepsilon>0$ small enough so that the circle
$\gamma_{\varepsilon}:=\{z:\ |z-\lambda|=\varepsilon\}$
lies in $\rho(\mathcal{K})$ and encloses no spectral point of $\mathcal{K}$ other than $\lambda$. Let $\mathcal{P}^{\gamma_{\varepsilon}}$ and $\mathcal{P}^{\gamma_{\varepsilon}}_{N}$ be the Riesz projections defined by $\gamma_{\varepsilon}$ (as in Definition 4.4). Then $\dim(\mathcal{P}^{\gamma_{\varepsilon}}\mathcal{F})=m$. By Corollary 4.9 applied to the contour $\gamma_{\varepsilon}$, for $N$ sufficiently large we have
$\dim(\mathcal{P}^{\gamma_{\varepsilon}}_{N}\mathcal{F})\geq m,$
hence $\mathcal{P}^{\gamma_{\varepsilon}}_{N}\neq 0$ and therefore $\sigma(\mathcal{K}_{N})\cap B(\lambda,\varepsilon)\neq\emptyset$. In particular, there exists $\lambda_{N}\in\sigma(\mathcal{K}_{N})\cap\Delta$ such that $|\lambda_{N}-\lambda|<\varepsilon$ for all large $N$.
Upper semicontinuity. Let $\overline{\Delta}$ be the closure of $\Delta$ and define the compact set
$\overline{\Delta}_{\varepsilon}:=\overline{\Delta}\setminus B(\lambda,\varepsilon).$
Since $\sigma(\mathcal{K})\cap\overline{\Delta}_{\varepsilon}=\emptyset$, we have $\overline{\Delta}_{\varepsilon}\subset\rho(\mathcal{K})$. Applying Lemma 4.6 to $U=\overline{\Delta}_{\varepsilon}$ yields $N(\varepsilon)$ such that for all $N>N(\varepsilon)$,
$\overline{\Delta}_{\varepsilon}\subset\rho(\mathcal{K}_{N}),$
and hence $\sigma(\mathcal{K}_{N})\cap\overline{\Delta}_{\varepsilon}=\emptyset$. Therefore every $\mu\in\sigma(\mathcal{K}_{N})\cap\Delta$ must satisfy $|\mu-\lambda|<\varepsilon$. Since $\varepsilon>0$ is arbitrary, every limit point of $\sigma(\mathcal{K}_{N})\cap\Delta$ equals $\lambda$.
Multiplicity preservation. Assume Assumption 4.5 holds, so $\dim(\mathcal{P}_{N}^{\lambda}\mathcal{F})=m$ for $N$ large enough. Then the finite-dimensional operator $\mathcal{P}_{N}^{\lambda}\mathcal{K}_{N}\mathcal{P}_{N}^{\lambda}$ acting on $\mathcal{P}_{N}^{\lambda}\mathcal{F}$ has dimension $m$, and its spectrum equals $\sigma(\mathcal{K}_{N})\cap\Delta$ (Riesz decomposition, cf. [4, Theorem 2.27]). Hence $\sigma(\mathcal{K}_{N})\cap\Delta$ consists of exactly $m$ eigenvalues, counting algebraic multiplicities. ∎
@@
Intuitively speaking, upper semicontinuity ensures that every limit of approximate spectral points is contained in the true spectrum, which prevents irrelevant points (i.e., spectral pollution) and lower semicontinuity ensures that every point of the true spectrum is approached by some approximate spectral points, which prevents loss of spectral points (i.e., spectral inclusion).
@@
The [4, Theorem 2.27] in the above theorem mainly says that one can construct a spectral projection via a contour integral, which decomposes the space into two invariant subspaces: one that contains exactly the isolated part of the spectrum and one that contains the remainder. Essentially, it provides a method to “split” or reduce $\mathcal{K}$ into simpler parts corresponding to these separated spectral regions.
@@
Now we want to extend the Hausdorff convergence result beyond isolated eigenvalues to a compact subset of the spectrum. In order to do so, we need to introduce a new concept called resolvent stability as follows.
@@
Let $\mathcal{K}$ be a bounded Koopman operator and $\{\mathcal{K}_{N}\}_{N\in\mathbb{N}}$ be finite-dimensional approximations that converge strongly to $\mathcal{K}$ and satisfy Assumption 4.2. Let $U\subset\sigma(\mathcal{K})$ be a non-empty compact subset enclosed by a positively oriented contour $\Gamma\subset\rho(\mathcal{K})$, and let $\Delta$ denote the interior of $\Gamma$. Assume $\sigma(\mathcal{K})\cap\Delta=U$ (equivalently, $\Delta\setminus U\subset\rho(\mathcal{K})$). Then
@@
Because $\Delta\setminus U\subset\rho(\mathcal{K})$ by assumption, every $z\in\Delta\setminus U$ lies in the resolvent set of $\mathcal{K}$. Assumption 4.2 therefore yields constants $N_{z}\in\mathbb{N}$ and $M_{z}>0$ such that for all $N>N_{z}$,
@@
where $\mathcal{R}_{N}(z)=(zI-\mathcal{K}_{N})^{-1}$.
Fix $z\in\Delta\setminus U$ and assume $\liminf_{N\to\infty}\operatorname{dist}\bigl(\sigma(\mathcal{K}_{N}),z\bigr)>0.$ Then there exist $\delta>0$ and $N_{0}$ such that $\operatorname{dist}(\sigma(\mathcal{K}_{N}),z)>\delta$ for all $N>N_{0}$. Consequently, for every $N>\max\{N_{0},N_{z}\}$ we have the uniform bound $\|\mathcal{R}_{N}(z)\|\leq M_{z}$. Hence the sequence $\{\mathcal{K}_{N}\}$ is resolvent-stable around $U$ in the sense of Definition 4.14 (take $N(U)=\Delta$).
Applying Mills’ theorem [25, Theorem 3.2], which requires strong convergence and resolvent stability around $U$, gives
which is the desired lower semicontinuity. ∎
@@
Upper semicontinuity in Theorem 4.16 is not satisfied unless we have a stronger condition that $\mathcal{K}_{N}$ uniformly converges to $\mathcal{K}$. Refer to [14, Theorem 3.1, Chapter 4] for more details.
We want to point out that Chatelin’s work in [4] primarily addresses the convergence of isolated eigenvalues which provides a comprehensive framework that ensures preservation of algebraic multiplicities during the approximation process. This is particularly valuable for numerical analysis as it guarantees that the number of approximate eigenvalues (counting multiplicity) matches that of the exact operator. In contrast, Mills’ work in [25] extends spectral convergence theory (i.e., lower semicontinuity) to compact subsets of the spectrum, which offers a much broader applicability beyond just isolated eigenvalues.
@@ 5 Error Bounds Analysis
In the last section, we established the convergence of the approximated spectrum $\sigma(\mathcal{K}_{N})$ to the true spectrum $\sigma(\mathcal{K})$. However, for practical application, it is essential to quantify the error of a computed solution. In this section, we will derive an error bound for the computed eigenvalues that is controlled by a residual quantity $\eta_{N}$, as defined later. This provides a practical way to assess the accuracy of the spectral information obtained via Galerkin methods such as EDMD [34] and our pseudo-resolvent approach.
@@
Let $\mathcal{P}_{N}^{\lambda}$ be the spectral projection associated with the eigenvalues of $\mathcal{K}_{N}$ inside the contour $\Gamma_{\lambda}$. We define the residual quantity $\eta_{N}$ as:
$\eta_{N}:=\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|.$
To support our main proof, we state a key lemma from spectral approximation theory in the following, which establishes the fundamental relationship between the error in the mean eigenvalue and the trace of an operator difference.
Lemma 5.3.
@@
Let $\lambda$ be an isolated eigenvalue of the Koopman operator $\mathcal{K}$ with finite algebraic multiplicity $m_{\lambda}$. Let $\Gamma_{\lambda}$ be a closed Jordan curve isolating $\lambda$ from the rest of $\sigma(\mathcal{K})$. Suppose that Assumption 4.2 and Assumption 4.5 hold. Then for $N$ sufficiently large, the error in the arithmetic mean of the approximate eigenvalues is bounded by the residual $\eta_{N}$ in Definition 5.2. Specifically, there exists a constant $C$ independent of $N$ such that
$|\lambda-\bar{\lambda}_{N}|\leq C\cdot\eta_{N}.$ (10)
@@
Let $\mathcal{M}_{N}^{\lambda}:=\mathcal{P}_{N}^{\lambda}\mathcal{F}$, which has dimension $m_{\lambda}$ for all large $N$ by Assumption 4.5. The eigenvalues of $\mathcal{K}_{N}$ inside $\Gamma_{\lambda}$ (counting algebraic multiplicity) are exactly the eigenvalues of the finite-dimensional operator $\mathcal{G}_{N}:=\mathcal{K}_{N}|_{\mathcal{M}_{N}^{\lambda}}$, hence
$\bar{\lambda}_{N}=\frac{1}{m_{\lambda}}\operatorname{tr}(\mathcal{G}_{N}).$
The desired a posteriori estimate is a direct consequence of [4, Theorem 6.15], which gives a constant $C$ independent of NN under stable and strongly stable convergence such that
$|\lambda-\bar{\lambda}_{N}|\leq C\,\|(\mathcal{K}-\mathcal{K}_{N})\mathcal{P}_{N}^{\lambda}\|.$
By Definition 5.2, the right-hand side equals $C\,\eta_{N}$. This completes the proof. ∎
Remark 5.5.
A numerical verification for this error bound is presented in Section 6.2.
@@ 6.1 Pendulum
In this experiment, data is generated by simulating a single undamped pendulum. From a physics perspective, this system is a conservative Hamiltonian system. Its trajectories preserve total energy, and the dynamics are characterized by the oscillations. Consequently, the true spectrum of the Koopman operator is expected to lie strictly on the unit circle, while the corresponding generator spectrum should lie on the imaginary axis.
@@
The simulation involves a single trajectory of 1500 time steps with a fixed sampling interval of 0.2 seconds. The system is initialized from a random state, where the angle $\theta$ is drawn uniformly from 0 to $2\pi$ and the angular velocity $\dot{\theta}$ from 5-5 to $5$. The underlying dynamics are governed by the nonlinear equation:
$\ddot{\theta}=-\frac{3g}{2l}\sin(\theta),$
@@
with gravitational acceleration $g=9.8$ and pendulum length $l=1.0$. The angular acceleration is computed at each step, and the state is updated via Euler’s method. For the dictionary, we employ a hybrid basis: a Fourier basis for the periodic angle $\theta$ (using $\exp\{in\theta\}$ for $n\in[-20,20]$) and a Hermite polynomial basis for the unbounded angular velocity $\dot{\theta}$ (up to order 30, with velocity rescaled by $1/\sqrt{2}$). These basis functions are multiplied to form the dictionary of observables used to approximate the resolvent and analyze the spectral properties.
@@
Figure 1 illustrates the behavior of the inverse resolvent norm, $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$, as the probe $z$ traverses circles with radii $\Gamma_{1.01}$, $\Gamma_{1.25}$, and $\Gamma_{1.5}$. Theoretically, as $z$ approaches the true spectrum (the unit circle), the resolvent norm should diverge, causing its inverse to vanish. This behavior is clearly observed: the value of $1/\|\mathbf{R}(z,\mathbf{K}_{N})\|$ decreases significantly as the radius approaches 1. Consequently, by refining the detection threshold, only grid points in close proximity to the unit circle are identified as spectral candidates, thereby enhancing the precision of the approximation.